OpenStudy (anonymous):

find the domain and range g(z)= 1 over square root 4-z square

OpenStudy (lgbasallote):

find the value of z that will make the equation undefined... sq root (4-z^2) = 0 4 - z^2 = 0 4 = z^2 z = +/- 2 so...your domain is all values of x except for 2 and -2...which can be written as...(-inf, -2)U(2, inf) as for the range...cross multiply. y(sq root 4-z^2) = 1 4 - z^2 = 1/y^2 z^2 = 4 - 1/y^2 z^2 = (4y^2 - 1 )/y^2 get the sq root z = sq root (2y +1)(2y-1)/y anyways...your range is everything below and equal to zero...so your range is (-inf, 0)U(0, inf)

OpenStudy (anonymous):

how you throw away the square root to find range?

OpenStudy (lgbasallote):

y(sq root 4-z^2) = 1 divide both sides by y sq root 4-z^2 = 1/y square both sides... 4-z^2 = 1/y^2

OpenStudy (anonymous):

hmmm \[\frac{1}{\sqrt{4-z^2}}\] need \[4-z^2>0\] \[(2-z)(2+z)>0\] i think this is \[(-2,2)\] as the domain

OpenStudy (anonymous):

it is a square root right?

OpenStudy (anonymous):

is z a real variable?

OpenStudy (anonymous):

ooooooooh range i see ok that would be \[(\frac{1}{2},\infty)\]

OpenStudy (anonymous):

reason as follows. this beast can get as large as possible by making the denominator small, that is make x close to 2 or -2 the denominator is largest when x is 0, and when x is 0 you get \[\frac{1}{\sqrt{4}}=\frac{1}{2}\] so that is the minimum value of the function

OpenStudy (anonymous):

so to be more precise, range is \[[\frac{1}{2},\infty)\] or \[\frac{1}{2}\leq y<\infty\]

OpenStudy (anonymous):

actly which answer the correct one??? im blur

OpenStudy (anonymous):

\[g(z)=\frac{1}{\sqrt{4-z^2}}\] Domain \[(-2,2)\] Range \[[\frac{1}{2},\infty)\]

OpenStudy (anonymous):

r u sure????

OpenStudy (anonymous):

why u say that x>0